OpenStudy (anonymous):

Consider a spherical cow that is consuming a great deal of hay. He is fattening himself up at a rate of 200 liters/day. How quickly is her radius growing when she can't fit through the circular stall door (which is 2 m wide)?

OpenStudy (anonymous):

@sourwing

OpenStudy (anonymous):

lol spherical cow

OpenStudy (anonymous):

teacher's crazy

OpenStudy (anonymous):

V = (4pi/3)r^3

OpenStudy (anonymous):

So, )4pi/3)(200)^3 right>?

OpenStudy (anonymous):

i got 3,200,000pi

OpenStudy (anonymous):

@SithsAndGiggles sourwing disappeared on me

OpenStudy (anonymous):

\[V=\frac{4}{3}\pi r^3~~\Rightarrow~~\frac{dV}{dt}=4\pi r^2\frac{dr}{dt}\] Given: \(\dfrac{dV}{dt}=200\) The cow won't be able to fit through the door when its diameter matches the width of the door: \(d=2r=2\text{m}~~\Rightarrow~~r=1\text{m}\). Keep in mind that \(1\text{L}=0.001\text{m}^3\). \[200\text{L}=0.2\text{m}^3=4\pi (1\text{m})^2\frac{dr}{dt}~~\Rightarrow~~\frac{dr}{dt}=\cdots\]

OpenStudy (anonymous):

so, the derivative of 4pi(1m)^2 right?

OpenStudy (anonymous):

No, we already found the derivatives we need. All you need to do is solve for \(\dfrac{dr}{dt}\).

OpenStudy (anonymous):

k i sec

OpenStudy (anonymous):

4pi?

OpenStudy (anonymous):

Not quite, check your algebra: \[\frac{dr}{dt}=\frac{0.2}{4\pi}\frac{\text{m}}{\text{s}}=\cdots\]

OpenStudy (anonymous):

It's a big number. i can't get fraction form. it's like 0.0159...

OpenStudy (anonymous):

You mean the answer is supposed to be a big number?

OpenStudy (anonymous):

0.015915494309?

OpenStudy (anonymous):

Yeah, that's what I got.

OpenStudy (anonymous):

so, there's no fraction form right?

OpenStudy (anonymous):

You could leave it as \(\dfrac{0.2}{4\pi}\), which is equal to \(\dfrac{1/5}{4\pi}=\dfrac{1}{20\pi}\)

OpenStudy (anonymous):

Yes! I see :D